OpenStudy (rainbow_rocks03):

PLZ help me I will give you a medal for the Best response.Which mixed number is equal to 7.6?

OpenStudy (rainbow_rocks03):

Is it 7 3/5?

OpenStudy (bee_see):

Yes. 

OpenStudy (anonymous):

Did you work out the problem? I got 7 3/5 also.

OpenStudy (igreen):

0.6 is the same as 6/10 Can you simplify that?

OpenStudy (igreen):

Oh, 7 3/5 is correct.. 0.6 = 6/10 = 3/5 So we have 7 3/5